OpenStudy (anonymous):

A charity fair raised $6,000 by selling 500 lottery tickets. There were two types of lottery tickets: A tickets cost $10 each, and B tickets cost $60 each. How many tickets of each type were sold?

OpenStudy (aripotta):

We have to set up a system of equations. So the total amount of tickets sold was 500. So A + B = 500. The total amount of money made was 6000. 10a + 60b = 6000 So our system is A + b = 50 10a + 60b = 6000

OpenStudy (anonymous):

ok... but what would be the next step im confused by math

OpenStudy (aripotta):

There are multiple ways to solve a system of equations. Let's use the elimination method. We want to eliminate a variable by combining the equations. But first, we need to make that possible. Let's work to eliminate the a variable. What can we do to the first equation to make it so the a It's eliminated when we combine?

OpenStudy (aripotta):

Is* not it's, sorry. I'm on a tablet

OpenStudy (anonymous):

its ok but i started plugging in answer choices with the formula and i got 10(480) + 60(20) = 4800 + 1200 = 6000

OpenStudy (anonymous):

so its 480 and 20

OpenStudy (aripotta):

Well I guess that's one way to get your answer lol. Do you still wanna work it out with me, though? So you know what to do on other problems?

OpenStudy (anonymous):

no its fine but thanks

OpenStudy (aripotta):

Alrighty. Best of luck~